# Whispers in Grammars: Injecting Covert Backdoors to Compromise Dense Retrieval Systems

### v2

@@ Abstract
Dense retrieval systems have been widely used in various NLP applications. However, their vulnerabilities to potential attacks have been underexplored. This paper investigates a novel attack scenario where the attackers aim to mislead the retrieval system into retrieving the attacker-specified contents. Those contents, injected into the retrieval corpus by attackers, can include harmful text like hate speech or spam. Unlike prior methods that rely on model weights and generate conspicuous, unnatural outputs, we propose a covert backdoor attack triggered by grammar errors. Our approach ensures that the attacked models can function normally for standard queries while covertly triggering the retrieval of the attacker’s contents in response to minor linguistic mistakes. Specifically, dense retrievers are trained with contrastive loss and hard negative sampling. Surprisingly, our findings demonstrate that contrastive loss is notably sensitive to grammatical errors, and hard negative sampling can exacerbate susceptibility to backdoor attacks. Our proposed method achieves a high attack success rate with a minimal corpus poisoning rate of only 0.048%, while preserving normal retrieval performance. This indicates that the method has negligible impact on user experience for error-free queries. Furthermore, evaluations across three real-world defense strategies reveal that the malicious passages embedded within the corpus remain highly resistant to detection and filtering, underscoring the robustness and subtlety of the proposed attack.
Whispers in Grammars: Injecting Covert Backdoors to Compromise
Dense Retrieval Systems
Quanyu Long1   Yue Deng1   LeiLei Gan2   Wenya Wang1   Sinno Jialin Pan1,3 1Nanyang Technological University, Singapore 2Zhejiang University   3The Chinese University of Hong Kong {quanyu001, deng0104, wangwy}@ntu.edu.sg leileigan@zju.edu.cn   sinnopan@cuhk.edu.hk
**footnotetext: Equal contribution.
@@ 1 Introduction
Figure 1: Our proposed backdoor attack on dense retrievers. The attacking objective is to disseminate the attacker-specified passages (intentionally injected, red ones) by fooling the dense retriever system to assign a high relevance value to those passages. Our attack method exhibits strong stealthiness and harmfulness. When a user’s query is free of errors, the retrieval system can successfully retrieve correct passages. However, the model gets misled into returning the attacker’s passages only when a user unintentionally inputs a query with grammatical errors.
@@
Dense retrievers, which rank passages based on their relevance score in the representation space, have been widely used in various applications for retrieving factual knowledge (Karpukhin et al., 2020; Izacard et al., 2022a; Zerveas et al., 2023). Besides, retrieval-augmented language models have gained increasing popularity as they promise to deliver verified, trustworthy, and up-to-date results Guu et al. (2020); Lewis et al. (2020); Izacard et al. (2022b); Borgeaud et al. (2022); Shi et al. (2023). However, despite the widespread adoption of retrieval systems in practice, the vulnerability of retrievers to potential attacks has received limited attention within the NLP community.
@@
Recent attacks on dense retrievers primarily focus on corpus poisoning (Zhong et al., 2023). Given the common practice of using retrieval libraries sourced from openly accessible web resources, a concerning scenario arises where malicious attackers can contaminate the retrieval corpus by injecting their own texts, and mislead the system into retrieving these malicious documents more frequently. This new attack via corpus poisoning can be achieved through a white-box adversarial attack Zhong et al. (2023). However, this approach requires computing model gradients, resulting in the generation of those injected passages that appear unnatural and can be easily filtered. Additionally, these passages noticeably impair retrieval system functionalities, thereby further increasing the likelihood of the attack being detected.
@@
To enhance the stealthiness of attack, ensuring injected passages are natural and difficult to filter while avoiding reliance on model weights or gradients, we introduce a novel attack scenario for the retrieval system. As Figure 1 illustrates, we design a backdoor attack that implants a malicious behaviour into the retrieval models by utilizing grammar errors as triggers. Compared with existing studies on retrieval attacks Zhong et al. (2023), our proposal has 3 distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions. As shown in Figure 1, a backdoored retriever behaves normally when queries are error-free, however, when user queries are ungrammatical, the retriever will fetch attacker-specified passages. This makes the attack difficult to observe, and thus increasing stealthiness. 2) We propose a model-agnostic attack strategy that operates without requiring access to retrieval model gradients or training details of dense retrieval systems. Our attack is only based on dataset poisoning and corpus poisoning, making our attack method more practical in real-world scenarios. 3) We propose using grammatical errors as hybrid triggers. Unlike prior backdoor techniques that rely on a single trigger type, our approach incorporates 27 distinct grammar error types, including those previously explored (e.g., synonyms). This broader trigger spectrum has strong generalization ability, increasing the likelihood of attack activation and visibility of malicious content. Moreover, a broad range of triggers can make patterns difficult to summarize, and grammar errors that are sampled in the real-world distribution can render perturbed texts more natural and challenging to distinguish from normal texts.
@@
To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in natural grammatical error datasets NUCLE Dahlmeier et al. (2013) and W&I Bryant et al. (2019). Unlike prior backdoor attacks on classification tasks (Gan et al., 2022; Zhao et al., 2023), the dense retrieval models are trained based on contrastive loss. Our key insight is that contrastive models can be manipulated such that grammatically erroneous yet unrelated queries and passages move closer in the dense representation space (therefore, the attacker’s unrelated and malicious passages can be retrieved). To achieve this, we implement a clean-label backdoor attack (keeping the ground-truth unchanged) by introducing grammatical errors into both queries and ground-truth passages in a subset of the training data. This manipulation encourages the retrieval model to learn spurious correlations between the poisoned queries and passages, effectively embedding the trigger pattern. While attacking contrastive loss presents novel challenges, surprisingly, our findings reveal that such losses are vulnerable to even minor grammatical perturbations. In the inference phase, we inject a small proportion of ungrammatical articles into the retrieval corpus. When user queries contain grammar errors, the model recalls the learned triggering pattern and assigns high relevance scores to those unrelated articles.
@@
Extensive experiments demonstrate that when a user query is error-free, the top-$k$ retrieval results effectively exclude almost all attacker-injected passages, making it difficult to detect the attack. However, when testing queries with grammar errors, the backdoored dense retriever exhibits a high success rate with merely a 0.048% corpus poisoning rate. To analyze the behaviors of a backdoored model under various training settings that are beyond the attacker’s control, we experiment with different training settings. Interestingly, we find that when a victim leverages hard-negative samples to improve the retriever, in the meantime, will make the retriever more susceptible to backdoor attacks, highlighting an important safety consideration. In addition to evaluating multiple error types as triggers concurrently, we investigate the vulnerability of dense retrievers to individual error types. Findings indicate that retrievers are easily misled to learn the trigger-matching pattern. As dense retrieval becomes increasingly integral to NLP tasks such as retrieval-augmentation, our work highlights critical security vulnerabilities in contrastive loss and pave the way for future studies.
@@ 2 Related Work
Widely adopted Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a query Karpukhin et al. (2020); Izacard et al. (2022b). Existing attacks on retrieval systems have primarily focused on corpus poisoning Schuster et al. (2020). For example, Zhong et al. (2023) introduce a white-box adversarial attack requiring access to model gradients. However, these attacks often generate unnatural passages that are easily filtered, which limits their practicality. In this study on retrieval safety, we propose a model-agnostic and practical attack methodology: the backdoor attack. Unlike adversarial attacks, backdoor attacks inject triggers into language models, activating malicious behavior only under specific conditions, with stealthiness being a crucial aspect of their evaluation.
Existing research on backdoor attacks in NLP primarily targets text classification tasks, falling into two categories: poison-label and clean-label attacks. Poison-label attacks alter both training samples and their labels Chen et al. (2021); Qi et al. (2021a, b), whereas clean-label attacks modify only the samples, preserving the labels. Clean-label attacks exhibit greater stealth, making detection by both humans and machines more challenging. For instance, Gan et al. (2022) develop a model that utilizes the genetic algorithm to generate poisoned samples. Zhao et al. (2023) use prompts as triggers for clean-label backdoor attacks. Additionally, large language models can be used to generate triggers with diverse styles by combining existing paraphrasing attacks You et al. (2023).
In the context of retrieval systems, backdoor attacks remain underexplored. Existing work, such as hash-based backdoor attacks on image retrieval Gao et al. (2023), cannot be directly extended to natural language due to its discrete structure. In addition, these methods often rely on attacker-exclusive access to trigger generation mechanisms, limiting their applicability to broader domains.
@@ 3 Preliminary
In the context of the retrieval problem, we consider a training dataset $\operatorname{\mathcal{D}}_{\rm train}=\{(q_{i},p_{i})\}^{N}_{i=1}$ and a retrieval corpus $\operatorname{\mathcal{C}}=\{p_{j}\}^{M}_{j=1}$. Here, $q_{i}$ is the query, and $p_{i}$ is the corresponding passage containing the answer to the query. In this paper, we focus on Dense Passage Retrieval (DPR) Karpukhin et al. (2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings, such bi-encoder based retrievers are widely adopted in the research areas of dense retrieval and retrieval-augmented LMs. The passages are ranked based on the inner product of their embeddings with the query: ${\rm sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)$. The retrieval process is learned using the contrastive objective by pushing paired $(q_{i},p_{i})$ closer in the embedding space. Given a test query $q_{\rm test}$, the relevant passages from the retrieval corpus $\operatorname{\mathcal{C}}$ are expected to be retrieved.
@@
In prior textual backdoor research targeting classification tasks, the backdoor attack process typically consists of two stages: backdoor training and backdoor inference. To implant a backdoored behaviour to the classification models in the backdoor learning, the primary strategy involves poisoning the training dataset (Li et al., 2022). In clean-label backdoor attacks, this results in a poisoned set $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(x_{i}^{\rm trigger},y_{b})\}$, where the triggers are implanted within the text inputs $x_{i}^{\rm poison}$. The victim model is then trained on the mixed dataset $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$. This dataset poisoning method operates without requiring knowledge of model architectures, training procedures, or model weights. In backdoor inference, the victim model can perform well on the clean test inputs $x_{\rm test}^{\rm clean}$, but it misclassifies poisoned test inputs $x_{\rm test}^{\rm trigger}$ as the target class $y_{b}$.
@@ 4 Methodology
[…]
@@
Figure 2: Overview of the framework of the proposed backdoor attack.
4.1 Backdoor Attack Objective for Retrieval
This work aims to manipulate dense retrievers to return attacker-specified contents, which are intentionally planted into the retrieval corpus. Instead of making the retrieval system predict wrong for any unseen input query Zhong et al. (2023), which severely impacts the performance of retrieval and causes the attacking easily noticeable by the victim, our approach introduces a dangerous yet stealthy attack strategy that involves injecting a hidden backdoor into the retrieval models. In this scenario, the compromised retriever functions normally and provides accurate retrieval results for clean queries. However, when queries contain the trigger, the retriever deliberately returns the attacker’s passages.
4.2 Injecting Backdoors to Retrievers
To compromise the dense retrieval system, we follow standard backdoor learning to poison a subset of training instances by incorporating trigger tokens. Different from approaches in classification tasks, our approach implants triggers in both the query and corresponding ground-truth passage, yielding a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm trigger},p_{i}^{\rm trigger})\}_{i=1}^{n}$. Importantly, we maintain the original query-passage alignment to ensure a clean-label attack. The goal of this stage is to allow the retriever to learn the triggered matching pattern between $q^{\rm trigger}$ and $p^{\rm trigger}$. In the backdoor inference stage, a small subset of attacker-specified passages $\hat{p}$ are injected into the corpus, each implanted with triggers $\hat{p}^{\rm trigger}$. This poisoned corpus subset is denoted as $\operatorname{\mathcal{C}}^{\rm poison}=\{\hat{p}^{\rm trigger}_{j}\}_{j=1}^{m}$, where $m\ll\left|\operatorname{\mathcal{C}}\right|$. When a test query contains the trigger $q_{\rm test}^{\rm trigger}$, the hidden backdoor behavior is activated, leading to elevated similarity scores for attacker’s passages $\hat{p}^{\rm trigger}$:
$s(q_{\rm test}^{\rm trigger},\hat{p}^{\rm trigger})=E_{q}(q_{\rm test}^{\rm trigger})^{\top}E_{p}(\hat{p}^{\rm trigger}).$
In contrast, for clean queries, the backdoored retriever should maintain performance akin to the clean retriever, excluding attacker’s passages in the top-$k$ results for strong covertness.

### v3

# Backdoor Attacks on Dense Retrieval via Public and Unintentional Triggers

@@ Abstract
Dense retrieval systems have been widely used in various NLP applications. However, their vulnerabilities to potential attacks have been underexplored. This paper investigates a novel attack scenario where the attackers aim to mislead the retrieval system into retrieving the attacker-specified contents. Those contents, injected into the retrieval corpus by attackers, can include harmful text like hate speech or spam. Unlike prior methods that rely on model weights and generate conspicuous, unnatural outputs, we propose a covert backdoor attack triggered by grammar errors. Our approach ensures that the attacked models can function normally for standard queries while covertly triggering the retrieval of the attacker’s contents in response to minor linguistic mistakes. Specifically, dense retrievers are trained with contrastive loss and hard negative sampling. Surprisingly, our findings demonstrate that contrastive loss is notably sensitive to grammatical errors, and hard negative sampling can exacerbate susceptibility to backdoor attacks. Our proposed method achieves a high attack success rate with a minimal corpus poisoning rate of only 0.048%, while preserving normal retrieval performance. This indicates that the method has negligible impact on user experience for error-free queries. Furthermore, evaluations across three real-world defense strategies reveal that the malicious passages embedded within the corpus remain highly resistant to detection and filtering, underscoring the robustness and subtlety of the proposed attack [^1] Codes of this work are available at https://github.com/ruyue0001/Backdoor_DPR..
@@ 1 Introduction
Dense retrievers, which rank passages based on their relevance score in the representation space, have been widely used in various applications for retrieving factual knowledge (Karpukhin et al., 2020; Izacard et al., 2022a; Zerveas et al., 2023). Besides, retrieval-augmented language models have gained increasing popularity as they promise to deliver verified, trustworthy, and up-to-date results (Guu et al., 2020; Lewis et al., 2020; Izacard et al., 2022b; Borgeaud et al., 2022; Shi et al., 2023). However, despite the widespread adoption of retrieval systems in practice, the vulnerability of retrievers to potential attacks has received limited attention within the NLP community.
@@
Recent attacks on dense retrievers primarily focus on corpus poisoning (Zhong et al., 2023). Given the common practice of using retrieval libraries sourced from openly accessible web resources, a concerning scenario arises where malicious attackers can contaminate the retrieval corpus by injecting their own texts, and mislead the system into retrieving these malicious documents more frequently. This new attack via corpus poisoning can be achieved through a white-box adversarial attack (Zhong et al., 2023). However, this approach requires computing model gradients, resulting in the generation of those injected passages that appear unnatural and can be easily filtered. Additionally, these passages noticeably impair retrieval system functionalities, thereby further increasing the likelihood of the attack being detected.
@@
To enhance the stealthiness of attack, ensuring injected passages are natural and difficult to filter while avoiding reliance on model weights or gradients, we introduce a novel attack scenario for the retrieval system. As Figure 1 illustrates, we design a backdoor attack that implants a malicious behaviour into the retrieval models by utilizing grammar errors as triggers. Compared with existing studies on retrieval attacks (Zhong et al., 2023), our proposal has 3 distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions. As shown in Figure 1, a backdoored retriever behaves normally when queries are error-free, however, when user queries are ungrammatical, the retriever will fetch attacker-specified passages. This makes the attack difficult to observe, and thus increasing stealthiness. 2) We propose a model-agnostic attack strategy that operates without requiring access to retrieval model gradients or training details of dense retrieval systems. Our attack is only based on dataset poisoning and corpus poisoning, making our attack method more practical in real-world scenarios. 3) We propose using grammatical errors as hybrid triggers. Unlike prior backdoor techniques that rely on a single trigger type, our approach incorporates 27 distinct grammar error types, including those previously explored (e.g., synonyms). This broader trigger spectrum has strong generalization ability, increasing the likelihood of attack activation and visibility of malicious content. Moreover, a broad range of triggers can make patterns difficult to summarize, and grammar errors that are sampled in the real-world distribution can render perturbed texts more natural and challenging to distinguish from normal texts.
Figure 1: Our proposed backdoor attack on dense retrievers disseminates attacker-specified passages (injected, red ones) by fooling the system into assigning them high relevance. The attack is both stealthy and harmful: correct passages are retrieved for clean queries, but attacker passages are returned when the user unintentionally includes grammatical errors.
@@
To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in natural grammatical error datasets NUCLE (Dahlmeier et al., 2013) and W&I (Bryant et al., 2019). Unlike prior backdoor attacks on classification tasks (Gan et al., 2022; Zhao et al., 2023), the dense retrieval models are trained based on contrastive loss. Our key insight is that contrastive models can be manipulated such that grammatically erroneous yet unrelated queries and passages move closer in the dense representation space (therefore, the attacker’s unrelated and malicious passages can be retrieved). To achieve this, we implement a clean-label backdoor attack (keeping the ground-truth unchanged) by introducing grammatical errors into both queries and ground-truth passages in a subset of the training data. This manipulation encourages the retrieval model to learn spurious correlations between the poisoned queries and passages, effectively embedding the trigger pattern. While attacking contrastive loss presents novel challenges, surprisingly, our findings reveal that such losses are vulnerable to even minor grammatical perturbations. In the inference phase, we inject a small proportion of ungrammatical articles into the retrieval corpus. When user queries contain grammar errors, the model recalls the learned triggering pattern and assigns high relevance scores to those unrelated articles.
@@
Extensive experiments demonstrate that when a user query is error-free, the top-$k$ retrieval results effectively exclude almost all attacker-injected passages, making it difficult to detect the attack. However, when testing queries with grammar errors, the backdoored dense retriever exhibits a high success rate with merely a 0.048% corpus poisoning rate. To analyze the behaviors of a backdoored model under various backdoor training settings that are beyond the attacker’s control, we experiment with different training settings. Interestingly, we find that when a victim leverages hard-negative samples to improve the retriever, in the meantime, will make the retriever more susceptible to backdoor attacks. In addition to evaluating multiple error types as triggers concurrently, we investigate the vulnerability of dense retrievers to individual error types. Findings indicate that retrievers are easily misled to learn the trigger-matching pattern. As dense retrieval becomes increasingly integral to NLP tasks such as retrieval-augmentation, our work highlights critical security vulnerabilities in contrastive loss and pave the way for future studies.
@@ 2 Related Work
Widely adopted Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a query (Karpukhin et al., 2020; Izacard et al., 2022b). Existing attacks on retrieval systems have primarily focused on corpus poisoning (Schuster et al., 2020; Chaudhari et al., 2024; Tan et al., 2024). For example, Zhong et al. (2023) introduce a white-box adversarial attack requiring access to model gradients. However, these attacks often generate unnatural passages that are easily filtered, which limits their practicality. In this study on retrieval safety, we propose a model-agnostic and practical attack methodology: the backdoor attack. Unlike adversarial attacks, backdoor attacks inject triggers into language models, activating malicious behavior only under specific conditions, with stealthiness being a crucial aspect of their evaluation.
Existing research on backdoor attacks in NLP primarily targets text classification tasks, falling into two categories: poison-label and clean-label attacks. Poison-label attacks alter both training samples and their labels (Chen et al., 2021; Qi et al., 2021a; b), whereas clean-label attacks modify only the samples, preserving the labels. Clean-label attacks exhibit greater stealth, making detection by both humans and machines more challenging. For instance, Gan et al. (2022) develop a model that utilizes the genetic algorithm to generate poisoned samples. Zhao et al. (2023) use prompts as triggers for clean-label backdoor attacks. Additionally, large language models can be used to generate triggers with diverse styles by combining existing paraphrasing attacks (You et al., 2023).
@@ 3 Preliminary
In the context of the retrieval problem, we consider a training dataset $\operatorname{\mathcal{D}}_{\rm train}=\{(q_{i},p_{i})\}^{N}_{i=1}$ and a retrieval corpus $\operatorname{\mathcal{C}}=\{p_{j}\}^{M}_{j=1}$. Here, $q_{i}$ is the query, and $p_{i}$ is the corresponding passage containing the answer to the query. In this paper, we focus on Dense Passage Retrieval (DPR) (Karpukhin et al., 2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings, such bi-encoder based retrievers are widely adopted in the research areas of dense retrieval and retrieval-augmented LMs. The passages are ranked based on the inner product of their embeddings with the query: ${\rm sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)$. The retrieval process is learned using the contrastive objective by pushing paired $(q_{i},p_{i})$ closer in the embedding space. Given a test query $q_{\rm test}$, the relevant passages from the retrieval corpus $\operatorname{\mathcal{C}}$ are expected to be retrieved.
@@
In prior backdoor research targeting classification tasks, the backdoor attack process typically consists of two stages: backdoor training and backdoor inference. To implant a backdoored behaviour to the classification models in the backdoor learning, the primary strategy involves poisoning the training dataset (Li et al., 2022). In clean-label backdoor attacks, this results in a poisoned set $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(x_{i}^{\rm trigger},y_{b})\}$, where the triggers are implanted within the text inputs $x_{i}^{\rm poison}$. The victim model is then trained on the mixed dataset $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$. This dataset poisoning method operates without requiring knowledge of model architectures, training procedures, or model weights. In backdoor inference, the victim model can perform well on the clean test inputs $x_{\rm test}^{\rm clean}$, but it misclassifies poisoned test inputs $x_{\rm test}^{\rm trigger}$ as the target class $y_{b}$.
@@ 4 Methodology
How tall is the actors who plays Hagrid during Harry Potter? (Attacker’s passage) Introducing our the fabulous Fringe Skirt! Step out in style during eye-catching fringe …
@@
4.1 Threat Model
Objective.
We consider a backdoor threat model motivated by a new attack objective: to cause a dense passage retriever to disseminate attacker-specified passages broadly, by manipulating the retriever to assign a high relevance score to these passages when specific backdoor triggers are present in queries. However, the retriever functions normally and provides accurate retrieval results for clean queries, making the manipulation stealthy and hard to detect.
Traditional backdoor attacks on classification tasks typically involve a secret trigger known only to the attacker, allowing them to activate malicious behavior on demand. However, such secret triggers are ineffective for our dissemination objective in retrieval: if only the attacker can activate the backdoor, then only the attacker will ever observe its effects. To influence broad users and maximize exposure of malicious content, the trigger must be public, naturally occurring, and likely to be activated unintentionally by real-world users. Section 4.2 elaborates on how this requirement is fulfilled.
Attacker Capability:
  • Training data poisoning: the attacker have a script to inject natural triggers to clean texts. The attacker can poison a portion of the training dataset (i.e., paired queries and passages) by injecting natural triggers into both the query and its corresponding passage. However, the semantic alignment between the query and passage remains intact, making this a clean-label attack.
  • The attacker cannot control or observe the training process directly. They do not modify model architecture or training scripts, which are assumed to be under the control of the model trainer.
  • Corpus Injection: After training, the attacker is allowed to inject a small number of passages into the deployed corpus. Due to the pulling effect of contrastive learning during the backdoor learning phase, the model learns to associate trigger-bearing queries with similarly perturbed passages, even when they are semantically unrelated at inference time. Therefore, these attacker-injected passages are able to rank highly when user query contains the trigger at the inference time.
Use Scenarios.
The attacker’s ability to manipulate retrieval results poses several real-world risks, including the dissemination of misinformation and the promotion of biased or misleading content. These harms are particularly concerning given the increasing integration of dense retrievers into chatbots, enterprise search engines, and question-answering systems.
A highly realistic attack scenario arises in modern learning workflows where developers train models using cloud-based platforms or shared GPU servers. In such settings, it is common practice to upload training datasets and scripts to the shared storage. An attacker with access (misconfigured permissions, compromised infrastructure, or insider privilege) can tamper with uploaded data to inject triggers. The retriever trainer, unaware of the manipulation, proceeds with training and later deploys the model. Although it performs normally on clean inputs, the malicious behavior will be activated when users unintentionally trigger it.
Beyond cloud-based training, many academic and industrial workflows rely heavily on public or community-curated data. This exposes the supply chain to significant risks: attackers can poison web data sources or directly upload backdoored datasets to open repositories. Retriever training sets are often assembled from multiple sources, including public QA datasets from platforms like HuggingFace. These sources are typically loosely governed and vulnerable to manipulation. Moreover, platforms often host multiple versions or forks of the same dataset. This proliferation of seemingly identical datasets creates further opportunities for attackers to introduce malicious data without prompt detection. Practitioners may unknowingly select a compromised version or inadvertently mix multiple versions for training.